### v1

# Automorphisms of Plane Curves defined from Chebychev polynomials

Abstract.
In this paper, we study the geometry and automorphism groups of the algebraic curves $\mathcal{C}_{d}$ defined by the equation $y^{d}=\varphi_{d}(x)$ over a field $k$ with $\operatorname{char}(k)\nmid 2d$, where $\varphi_{d}(x)$ is the Chebyshev polynomial of degree $d$. We classify the total inflection points of $\mathcal{C}_{d}$, correcting and extending previous work on this. Additionally, we determine the automorphism groups of $\mathcal{C}_{d}$ in several cases, namely for $d=4$, and for any $d$ such that $12d=q+$ 1 or $14d=q+$ 1 for an arbitrary power $q$ of the prime $p=\operatorname{char}(k)$. As an application, we use our results to show that certain maximal curves (over finite fields) of the same genus are not isomorphic.
1. Introduction and preliminaries
For a field $k$ and an integer $d\geq 1$, the $d$-th Chebyshev polynomial $\varphi_{d}(X)$ is the unique monic polynomial such that in the ring of Laurent polynomials $k[t,t^{-1}]$ the equality
$\varphi_{d}(t+t^{-1})=t^{d}+t^{-d}$
holds. Equivalently, considering $x\mapsto\varphi_{d}(x)$ as a morphism ${\mathbb{P}}^{1}\to{\mathbb{P}}^{1}$ defined over $k$, the polynomial is characterized by the condition that the diagram
$\begin{array}[]{lcl}{{\mathbb{P}}}^{1}&\stackrel{{\scriptstyle r\mapsto r^{d}}}{{\longrightarrow}}&\mathbb{P}^{1}\\ {\Big{\downarrow}}{\scriptstyle r\mapsto r+r^{-1}}&&{\Big{\downarrow}}{\scriptstyle s\mapsto s+s^{-1}}\\ \mathbb{P}^{1}&\stackrel{{\scriptstyle t\mapsto\varphi_{d}(t)}}{{\longrightarrow}}&\mathbb{P}^{1}\end{array}$
commutes. From the diagram (see also [12] and [1, Remark 6.2]) it is readily deduced that $\varphi_{d}(X)\in k[X]$ is separable, except in the case that $\operatorname{char}(k)$ divides $2d$.
Throughout this paper it is assumed that $\operatorname{char}(k)\nmid 2d$. Define the curve ${\mathcal{C}}_{d}$ over $k$ as the closure in ${\mathbb{P}}^{2}$ of the affine curve given by
$y^{d}=\varphi_{d}(x).$
Since $\varphi_{d}(X)\in k[X]$ is separable, ${\mathcal{C}}_{d}$ is a smooth curve. Our main goal is to describe the group $\operatorname{Aut}({\mathcal{C}}_{d})$ consisting of the automorphisms of the curve ${\mathcal{C}}_{d}$ defined over a fixed algebraic closure $\overline{k}$ of $k$. We restrict to the case $d\geq 4$, or equivalently, the genus $g({\mathcal{C}}_{d})=(d-1)(d-2)/2\geq 2$, to ensure that $\operatorname{Aut}({\mathcal{C}}_{d})$ is a finite group. A classical result [10] of B. Segre (see also [5, Theorem 11.29]) yields that automorphisms of any smooth curve in ${\mathbb{P}}^{2}$ extend to linear automorphisms of ${\mathbb{P}}^{2}$, hence $\operatorname{Aut}({\mathcal{C}}_{d})$ defines a finite subgroup of $\operatorname{PGL}(3,\overline{k})$.
Our main results concern the classification of total inflection points of $\mathcal{C}_{d}$ and a detailed description of $\operatorname{Aut}({\mathcal{C}}_{d})$ in several key cases. Specifically, we prove the following:
  1. (I)
    The total inflection points of $\mathcal{C}_{d}$ are precisely the $d$ points in $\mathcal{Z}(y)\cap\mathcal{C}_{d}$, except in the case where $\operatorname{char}(k)=p\geq 3$ and $d=\frac{p^{m}+1}{2}$ for some $m\geq 1$. In this exceptional case there are $3d$ total inflection points. This result corrects and extends [1, Theorem 6.3] (see Proposition 2.1).
  2. (II)
    If $12d=q+$ 1 for some power $q$ of $p$, which is the situation described in (I), then $\mathcal{C}_{d}$ is isomorphic to a Fermat curve of degree $d$ and its automorphism group is $(\mathbb{Z}/d\mathbb{Z}\times\mathbb{Z}/d\mathbb{Z})\rtimes S_{3}$ (see Propositions 3.1 and 3.3).
  3. (III)
    If $d=4$, we distinguish the following cases:
    • For $p=5$, the automorphism group of $\mathcal{C}_{d}$ is isomorphic to $\mathbb{Z}/4\mathbb{Z}\times A_{4}$ (see Remark 3.11).
    • For $p=7$, the curve is isomorphic to the Fermat curve, and the automorphism group of $\mathcal{C}_{d}$ is isomorphic to $(\mathbb{Z}/d\mathbb{Z}\times\mathbb{Z}/d\mathbb{Z})\rtimes S_{3}$ (see Propositions 3.1 and 3.3).
    • For $p\neq 2,5,7$, the automorphism group of $\mathcal{C}_{d}$ is isomorphic to $\mathbb{Z}/4\mathbb{Z}\rtimes(\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z})$ (see Proposition 3.10).
  4. (IV)
    If $14d=q+$ 1 for some power $q$ of $p$, then the automorphism group of $\mathcal{C}_{d}$ is $\mathbb{Z}/d\mathbb{Z}\rtimes S_{3}$ (see Theorem 3.17 and Proposition 3.18).
  5. (V)
    We show that for many integers $m$ and $n$ and finite fields ${\mathbb{F}}_{q^{2}}$, the maximal curves defined by $y^{n}=\varphi_{m}(x)$ and $y^{n}=x^{m}+1$ are not isomorphic despite having the same genus. This result provides a corrected proof of [1, Corollary 6.4] and also generalizes it (see Propositions 4.1, 4.2, and 4.3).
The organization of this paper is as follows. In Section 2, we analyze the total inflection points of $\mathcal{C}_{d}$ and establish the correction to [1]. Section 3 is dedicated to the classification of the automorphism groups in the cases $12d=q+$ 1 and $14d=q+$ 1, respectively. Finally, Section 4 discusses applications to maximal curves over finite fields.
2. Total inflection points
Given a smooth curve ${\mathcal{C}}\subset{\mathbb{P}}^{2}$ of degree $d$, a total inflection point of ${\mathcal{C}}$ is a point $P\in{\mathcal{C}}$ such that the associated tangent line to ${\mathcal{C}}$ meets ${\mathcal{C}}$ in $P$ with multiplicity $d$.
Since automorphisms of any such smooth plane curve ${\mathcal{C}}$ are given by linear maps on ${\mathbb{P}}^{2}$, they act as permutations on the set of total inflection points of ${\mathcal{C}}$. We will compute these points in the case of ${\mathcal{C}}_{d}$, assuming $d\geq 3$ and $\operatorname{char}(k)\nmid 2d$. The result will help us in our ultimate goal, namely obtaining information on $\operatorname{Aut}({\mathcal{C}}_{d})$. The computation extends the proof of [1, Theorem 6.3], where the special case $p:=\operatorname{char}(k)>2$ and $\overline{-1}\in\langle\,\overline{p}\,\rangle\subseteq(\mathbb{Z}/2d\mathbb{Z})^{\times}$ is considered. Moreover, our computation corrects some claims on p. 17 of [1]: for example, the reasoning presented there leads to the erroneous conclusion that ${\mathcal{C}}_{d}$ has either $d$ or $2d$ total inflection points, contradicting their assertion in [1, Theorem 6.3].
Proposition 2.1.
Let $d\in\mathbb{Z}_{\geq 4}$. The only total inflection points on the curve ${\mathcal{C}}_{d}\subset{\mathbb{P}}^{2}$ defined as the closure of the affine equation $y^{d}=\varphi_{d}(x)$ over a field $k$ with $\operatorname{char}(k)\nmid 2d$, are the $d$ points in ${\mathcal{Z}}(y)\cap{\mathcal{C}}_{d}$, except in the following situation (where there are $3d$ such points, namely the ones in the intersection $\mathcal{Z}(y\cdot(x+2)\cdot(x-2))\cap{\mathcal{C}}_{d}$):
  1. $\operatorname{char}(k)=p\geq 3$ and $d=(p^{m}+1)/2$ for any $m\in\mathbb{Z}_{\geq 1}$ such that $d\geq 4$.
Proof.
The curve ${\mathcal{C}}_{d}$ is defined by the homogeneous equation
$y^{d}=z^{d}\cdot\varphi_{d}(x/z)=:\Phi_{d}(x,z).$
We consider all possibilities for a line $\ell\subset{\mathbb{P}}^{2}$ with #$\#(\ell\cap{\mathcal{C}}_{d})=1$.
(1) The line at infinity ($z=0$) leads to $\ell\cap{\mathcal{C}}_{d}$ given by $y^{d}=x^{d}$, hence $d$ distinct intersection points.
(2) A line $\ell\colon x=az$ (for fixed $a$) results in an intersection given by $y^{d}=z^{d}\cdot\varphi_{d}(a)$. Hence this yields a single intersection point precisely when $a$ is a zero of $\varphi_{d}(X)$. The corresponding $d$ total hyperflexes are $(a:0:1)$, with tangent line $x=az$.
(3) A line $\ell\colon y=az$ leads to the affine equation $\varphi_{d}(x)=a^{d}$. From the diagram on page 1 it is clear that $t\mapsto\varphi_{d}(t)$ has, over points in ${\mathbb{P}}^{1}\setminus\{\infty\}$, only ramification of index $2$. Hence such a line $\ell$ intersects ${\mathcal{C}}_{d}$ in at least $d/2$ points.
(4) What remains, is the case $\ell\colon y=ax+bz$ where $a\neq 0$. This gives #$\#(\ell\cap{\mathcal{C}}_{d})=1$ precisely in the situation
$(ax+bz)^{d}-\Phi_{d}(x,z)=(c_{1}x+c_{2}z)^{d}.$
Here an equality $a^{d}-1=c_{1}^{d}$ is obtained by considering the coefficient of $x^{d}$, while the coefficient of $z^{d}$ results in $b^{d}-\varphi_{d}(0)=c_{2}^{d}$.
We claim that $c_{1}\neq 0$. Indeed, the assumption $c_{1}=0$ implies that $\varphi_{d}(x)=(ax+b)^{d}-c_{2}^{d}$. Now observe that the two maps $t\mapsto\varphi_{d}(t)$ and $t\mapsto(at+b)^{d}-c_{2}^{d}$ are clearly distinct (the first has only $\infty$ as a ramification point with index $d$ while the other is totally ramified over both $\infty$ and $-c_{2}^{d}$). As a consequence,
$a^{d}\left(x+\frac{b}{a}\right)^{d}-\varphi_{d}(x)=(a^{d}-1)(x+c_{2})^{d}.$
Writing
$\alpha:=a^{d},\;\beta:=b/a,\;\gamma:=c_{2},\;\text{and}\;x:=t+\frac{1}{t},$
one obtains
$\alpha\left(t+\beta+\frac{1}{t}\right)^{d}-t^{d}-t^{-d}=(\alpha-1)\left(t+\gamma+\frac{1}{t}\right)^{d}.$
Comparing coefficients results in the following system of equations.
$\begin{array}[]{lrclr}t^{d-1}\colon&\alpha\beta d&=&(\alpha\!-\!1)\gamma d&(1)\\ t^{d-2}\colon&\alpha\beta^{2}\binom{d}{2}+\alpha d&=&(\alpha\!-\!1)\left(\gamma^{2}\binom{d}{2}+d\right)&(2)\\ t^{d-3}\colon&\alpha\beta^{3}\binom{d}{3}+2\alpha\beta\binom{d}{2}&=&(\alpha\!-\!1)\left(\gamma^{3}\binom{d}{3}+2\gamma\binom{d}{2}\right)&(3)\\ t^{d-4}\colon&\alpha\!\left(\beta^{4}\binom{d}{4}+\beta^{2}d\binom{d-1}{2}+\binom{d}{2}\right)&=&(\alpha\!-\!1)\left(\gamma^{4}\binom{d}{4}+\gamma^{2}d\binom{d-1}{2}+\binom{d}{2}\right)&(4)\end{array}$
The remainder of the calculation distinguishes various possibilities for $d\in k$.
  • Assume that $d=1$ in the field $k$. The equality (2) results in the contradiction $\alpha=\alpha-1$.
  • Next, consider the case $d=2$ in $k$, with $\operatorname{char}(k)\neq 3$. Now equality (4) yields the contradiction $\alpha=\alpha-1$.
  • Assume $d-1\neq 0$ and $d-2\neq 0$ in $k$ (hence in particular $\operatorname{char}(k)\neq 3$). Then (1) and (3) yield the system
    $\left\{\begin{array}[]{rcl}\alpha\beta&=&(\alpha-1)\gamma\\ \alpha\beta^{3}(d-2)+6\alpha\beta&=&(\alpha-1)\gamma^{3}(d-2)+6(\alpha-1)\gamma,\end{array}\right.$
    or equivalently (add a suitable multiple of the first equation to the second and multiply by $(d-2)^{-1}$), one obtains
    $\left\{\begin{array}[]{rcl}\alpha\beta&=&(\alpha-1)\gamma\\ \alpha\beta^{3}&=&(\alpha-1)\gamma^{3}.\end{array}\right.$
    This means $\alpha\beta^{3}=(\alpha-1)\gamma\cdot\gamma^{2}=\alpha\beta\gamma^{2}$, so since $\alpha\neq 0$ we are in one of the cases
    $\beta=0\;\;\text{or}\;\;\beta=\gamma\;\;\text{or}\;\;\beta=-\gamma.$
    If $\beta=0$ then (1) and (2) result in the contradiction $\alpha=\alpha-1$.
    If $\beta=\gamma$ (which has to be nonzero by what we just showed), then (1) yields the same contradiction $\alpha=\alpha-1$.
    The conclusion is that $\beta=-\gamma$, and this is nonzero. Hence (1) implies $\alpha=2^{-1}$. Substituting these values of $\alpha,\beta$ in (1)-(4) leads to the system
    $\left\{\begin{array}[]{l}\gamma^{2}(d-1)=-2,\\ \gamma^{4}(d-2)(d-3)+\gamma^{2}\cdot 12(d-2)=-12.\end{array}\right.$
    So $\gamma^{2}=2/(1-d)$ and
    $\frac{4}{(d-1)^{2}}(d-2)(d-3)-\frac{24}{d-1}(d-2)+12=0,$
    resulting in $d\in\{3,2^{-1}\}\subset k$. This is not possible if $\operatorname{char}(k)=0$, since we assume $d\in\mathbb{Z}_{\geq 4}$. For $p=\operatorname{char}(k)\geq 5$, two situations remain:
    (*) if $d\equiv 3\bmod p$, one obtains
    $\left(t-i+\frac{1}{t}\right)^{d}+\left(t+i+\frac{1}{t}\right)^{d}=2t^{d}+2t^{-d}$
    with $i$ any primitive $4$-the root of unity. Write $d=3+qn$ with $q$ a power of $p$ and $\gcd(n,p)=1$. Comparing coefficients of $t^{qn-q+2}$ results in ±$\pm 6n\equiv 0\bmod p$, a contradiction. Hence this case cannot occur.
    (*) If $2d\equiv 1\bmod p$, one obtains $\gamma^{2}=4$ and
    $\left(t-2+\frac{1}{t}\right)^{d}+\left(t+2+\frac{1}{t}\right)^{d}=2t^{d}+2t^{-d}.$
    Taking $t=s^{2}$, this gives
    $\left(s-\frac{1}{s}\right)^{2d}+\left(s+\frac{1}{s}\right)^{2d}=2s^{2d}+2s^{-2d}.$
    Now write $2d=1+qn$ with $q$ a power of $p$ and $\gcd(n,p)=1$. If $n=1$, the given equality indeed holds in ${\mathbb{F}}_{p}[s,s^{-1}]$. For $n>1$, considering the coefficient of $s^{2d-2q-2}$ leads to the contradiction $2n\equiv 0\bmod p$.
  • The final case to consider, is $\operatorname{char}(k)=3$ and $d\equiv 2\bmod 3$. Writing $d=2+3m$, this means $\binom{d}{3}\equiv m\bmod 3$ and $\binom{d}{4}\equiv-m\bmod 3$. The system of equations derived on p. 2 then takes the form
    $\begin{array}[]{lrclcr}t^{d-1}\colon&2\alpha\beta&=&2(\alpha\!-\!1)\gamma&&(1)\\ t^{d-2}\colon&\alpha\beta^{2}+2\alpha&=&(\alpha\!-\!1)\left(\gamma^{2}+2\right)&&(2)\\ t^{d-3}\colon&m\alpha\beta^{3}+2\alpha\beta&=&(\alpha\!-\!1)\left(m\gamma^{3}+2\gamma\right)&&(3)\\ t^{d-4}\colon&\alpha\!\left(-m\beta^{4}+1\right)&=&(\alpha\!-\!1)\left(-m\gamma^{4}+1\right)&&(4)\end{array}$
    Combining (1) and (3) results, as in the previous case, in the equalities $\beta=-\gamma$ and $\alpha=2^{-1}=2$. Then (2), (4) reduce to the system
    $\left\{\begin{array}[]{c}2\gamma^{2}+1=\gamma^{2}+2,\\ m\gamma^{4}+2=-m\gamma^{4}+1\end{array}\right.$
    which is equivalent to $\gamma=\pm 2$ and $m=1\in k$. So in integers one obtains $d=2+3(1+3n)\equiv 5\bmod 9$, and in $k[t,t^{-1}]$ the equality
    $\left(t+2+\frac{1}{t}\right)^{d}+\left(t-2+\frac{1}{t}\right)^{d}=2t^{d}+2t^{-d}$
    holds. Writing $t=s^{2}$ and $2d=1+q\ell$ with $q\geq 9$ a power of $3$ and $3\nmid\ell$, this takes the form
    $(s+s^{-1})(s^{q}+s^{-q})^{\ell}+(s-s^{-1})(s^{q}-s^{-q})^{\ell}=2s^{2d}+2s^{-2d}.$
    In case $\ell=1$ (in $\mathbb{Z}$) this identity of Laurent polynomials indeed holds. In the remaining case $\ell\geq 2$, the coefficient of $s^{q\ell-2q-1}$ in the left-hand-side equals $2\ell\neq 0$ in $k$, hence the equality is not possible.
This completes the proof; the calculations show that indeed the total inflection points are the ones described in the statement of the proposition. ∎
Remark 2.2.
Note that in Proposition 2.1, in any case where only $d$ total inflection points exist, these points are all on a unique line (namely, on the line given by $y=0$). Similarly, in any case with $3d$ total inflection points ($d\geq 4$), there is a unique set of three lines such that their union intersected with ${\mathcal{C}}_{d}$ equals this set of points (namely, the lines defined by $y=0$ and $x=\pm 2$).
3. Automorphism groups
As a first step towards understanding the group of automorphisms of a plane curve ${\mathcal{C}}_{d}$ corresponding to $y^{d}=\varphi_{d}(x)$, we distinguish according to the number of total inflection points as given by Proposition 2.1.
Proposition 3.1.
Let $\mathcal{C}_{d}$ be the curve defined by
$\mathcal{C}_{d}:y^{d}=\varphi_{d}(x)$
over a field of characteristic $p>0$. If $12d=q+$ 1 with $q$ a power of $p$, then $\mathcal{C}_{d}$ is isomorphic over $\mathbb{F}_{q^{2}}$ to the Fermat curve
$\mathcal{F}:v^{d}=u^{d}+1.$
Proof.
Consider the map $\mathcal{F}\to{\mathbb{P}}^{2}$ given in affine coordinates by
$(u,v)\mapsto\left(x:=\frac{2u+2}{u-1}\,,\,y:=\frac{av}{u-1}\right)$
where $a^{d}=2$ and so $a\in{\mathbb{F}}_{q^{2}}$. Since $x,y$ generate the function field ${\mathbb{F}}_{q^{2}}(\mathcal{F})$, this map defines a birational map to its image. We will show that its image is the curve $\mathcal{C}_{d}$, in other words we show
$2v^{d}=(u-1)^{d}y^{d}=(u-1)^{d}\varphi_{d}(\frac{2u+2}{u-1})$
in case $12d=q+$ 1. Equivalently, we claim that for $d=(q+1)/2$, in the polynomial ring ${\mathbb{F}}_{p}[u]$ the equality $(u-1)^{d}\varphi_{d}(\frac{2u+2}{u-1})=2u^{d}+2$ holds. Indeed, write $u=s^{2}$. In $\mathbb{F}_{p}(s)\supset{\mathbb{F}}_{p}(u)$ one observes
$\frac{2u+2}{u-1}=\frac{s+1}{s-1}+\frac{s-1}{s+1}.$
From this, we obtain
$(u-1)^{d}\varphi_{d}\left(\frac{2u+2}{u-1}\right)=(s+1)^{d}(s-1)^{d}\left(\left(\frac{s+1}{s-1}\right)^{d}+\left(\frac{s-1}{s+1}\right)^{d}\right).$
Since $2d=1+q$, this equals
$(s^{q}+1)(s+1)+(s^{q}-1)(s-1)=2u^{d}+2.$
Hence we described an explicit birational map from $\mathcal{F}$ to ${\mathcal{C}}_{d}$. Since both curves are smooth, the result follows. ∎
Remark 3.2.
For $12d=q+$ 1, it is known that the curve $y^{d}=\varphi_{d}(x)$ is maximal over $\mathbb{F}_{q^{2}}$. Thus, by [2, Theorem 1.1], the curve is isomorphic to the maximal Fermat curve of degree $d=(q+1)/2$. Here we provided a direct proof of this result.
The explicit isomorphism obtained here can in fact be obtained using Proposition 2.1. Namely, it is well known that in terms of homogeneous coordinates $u,v,w$ describing the Fermat curve $\mathcal{F}\colon v^{d}=u^{d}+w^{d}$, the total inflection points of $\mathcal{F}$ are the $3d$ points in $\mathcal{F}\cup\mathcal{Z}(uvw)$. Moreover, $\operatorname{Aut}(\mathcal{F})$ permutes the three lines $\mathcal{Z}(u),\mathcal{Z}(v),\mathcal{Z}(w)$. The map used in the proof of Proposition 3.1 yields one of the possible ways to associate these lines to the analogous lines for ${\mathcal{C}}_{d}$ via a linear map.
Proposition 3.3.
Consider the curve $\mathcal{C}_{d}$ given by $y^{d}=\varphi_{d}(x)$ defined over a field of characteristic $p>0$. If $12d=q+$ 1 with $q$ a power of $p$, then the automorphism group $G$ of the curve $\mathcal{C}_{d}$ has order $6d^{2}$ and has a normal abelian subgroup $H\cong\mathbb{Z}/d\mathbb{Z}\times\mathbb{Z}/d\mathbb{Z}$ of order $d^{2}$ such that $G/H\cong S_{3}.$
Proof.
From Proposition 3.1, we know the curve $\mathcal{C}_{d}$ is isomorphic to the Fermat curve of degree $d$ and the result follows from [9] and [11]. ∎
This completes the determination of $\text{Aut}(\mathcal{C}_{d})$ whenever this curve has $3d$ total inflection points. From now on we consider the remaining case.
Lemma 3.4.
Suppose ${\mathcal{C}}_{d}$ contains precisely $d$ total inflection points. Extending automorphisms of ${\mathcal{C}}_{d}$ to linear automorphisms of $\mathbb{P}^{2}$, these restrict to automorphisms of the line $\mathcal{Z}(y)\cong\mathbb{P}^{1}$, and this defines a group homomorphism
$\textrm{\rm Aut}({\mathcal{C}}_{d})\longrightarrow\textrm{\rm Aut}(\mathbb{P}^{1}).$
Its kernel consists of the automorphisms $(x,y)\mapsto(x,\zeta y)$ with $\zeta$ ranging over the $d$-th roots of unity. Its image is the group of Möbius transformations $x\mapsto\frac{ax+b}{cx+\delta}$ that permute the set of zeros of $\varphi_{d}(x)$.
Proof.
The set of total inflection points of ${\mathcal{C}}_{d}$ is preserved under any automorphism (since automorphisms are given by linear maps. The assumption implies that $y=0$ defines the unique line containing all these inflection points, hence any automorphism restricts to a linear automorphism of this line, i.e., to a Möbius transformation. Since elements of $\text{\rm Aut}({\mathcal{C}}_{d})$ permute the total inflection points, these Möbius transformations permute the zeros of $\varphi_{d}(x)$. Vice versa, if $x\mapsto\frac{ax+b}{cx+\delta}$ permutes the zeros of $\varphi_{d}(x)$, then
$\varphi_{d}(\frac{ax+b}{cx+\delta})=e(cx+\delta)^{-d}\varphi_{d}(x)$
for some nonzero constant $e$. As a consequence, the given Möbius transformation is the image of the element
$(x,y)\mapsto\left(\frac{ax+b}{cx+\delta}\,,\,\sqrt[d]{e}\frac{y}{cx+\delta}\right)$
in $\text{\rm Aut}({\mathcal{C}}_{d})$.
If an automorphism is in the kernel, i.e., restricts to the identity map $x\mapsto x$, then indeed it is of the asserted form. ∎
Remark 3.5.
Denote by $G$ the automorphism group of ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$, and let the subgroup $N$ be given by
$N=\{(x,y)\mapsto(x,\zeta y)\mid\zeta^{d}=1\}.$
Thus, Lemma 3.4 asserts that if ${\mathcal{C}}_{d}$ contains precisely $d$ total inflection points, then $N$ is a normal subgroup, and $G/N$ is isomorphic to a finite subgroup of $\textrm{\rm Aut}(\mathbb{P}^{1}).$
It may be interesting to observe that Lemma 3.4 and the above proof generalize to any smooth plane curve ${\mathcal{D}}\colon y^{d}=g(x)$ where $g(x)$ is a separable polynomial of degree $d\geq 4$. Namely, the $d$ zeros $\alpha$ of $g(x)$ yield total inflection points $(\alpha,0)\in{\mathcal{D}}$. Hence if ${\mathcal{D}}$ contains precisely $d$ total inflection points, then the (automatically linear) automorphisms of ${\mathcal{D}}$ restrict to ${\mathcal{Z}}(y)\cong{\mathbb{P}}^{1}$ and one proceeds as in the proof of Lemma 3.4.
We will use some well known information regarding coefficients of the polynomials $\varphi_{d}(x)$. To verify them is straightforward; in fact already Dickson’s 1893 thesis published as [4] presents this. One writes
$\varphi_{d}(x)=\sum_{j=0}^{\lfloor d/2\rfloor}c_{j}x^{d-2j},$
then the coefficients $c_{j}$ are (the reductions modulo $p$ of) the integers
[…]
In particular, the coefficient of $x^{d-2}$ is $c_{1}=-d$ and that of $x^{d-4}$ is $c_{2}=d(d-3)/2$, and
$\begin{array}[]{llcc}\text{if}\;d\;\text{is odd:}&c_{(d-1)/2}=(-1)^{(d-1)/2}d,&&\\ \text{if}\;d\;\text{is even:}&c_{d/2}=(-1)^{d/2}\cdot 2&\text{and}&c_{(d-2)/2}=(-1)^{(d-2)/2}d^{2}/4.\end{array}$
The following lemma describes, under a mild condition on $d$, the automorphisms of ${\mathcal{C}}_{d}$ of a specific form.
Proposition 3.6.
Suppose that either $d$ is even, or $d$ is odd and $\gcd(p,d-1)=1$. The automorphisms of the curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ of the form
$(x,y)\mapsto(\frac{x}{ax+b},\frac{y}{ax+b})$
are only the identity and those given by
$(x,y)\mapsto(-x,\pm y).$
Proof.
Note that a necessary condition for a map of the given form to be birational, is that $b\neq 0$, since $a=\frac{x}{ax}$ and $\frac{y}{ax}$ do not generate the function field of ${\mathcal{C}}_{d}$. The existence of an automorphism of the stated form is equivalent to the equality of polynomials
$\varphi_{d}(x)=(ax+b)^{d}\varphi_{d}\left(\frac{x}{ax+b}\right).$
If $d$ is even, considering the coefficient of $x$, one obtains
±$\pm 2dab^{d-1}=0.$
This implies that $a=0$. Now comparing the terms of degree $d-2$ shows $-d=-db^{2}$, leading to $b=\pm 1$.
If $d$ is odd, considering the coefficient of $x^{2}$ yields
$d(d-1)ab^{d-2}=0,$
which implies that $a=0$. Again using the coefficient of $x^{d-2}$ one concludes $b=\pm 1$. ∎
The following characterization of subgroups of projective linear groups is given in [13, Theorem 3]:
Theorem 3.7.
The group $\mathrm{PGL}(2,p^{m})$ has only the following subgroups:
  1. (1)
    Elementary abelian $p$-groups of order $p^{f}$ with $f\mid m$;
  2. (2)
    Cyclic groups of order $n$ with $n\mid(p^{m}\pm 1)$;
  3. (3)
    Dihedral groups $D_{n}$ of order $2n$ with $n\mid(p^{m}\pm 1)$;
  4. (4)
    $A_{4}$, for $p>2$ or $p=2$ and $m\equiv 0\pmod{2}$;
  5. (5)
    $S_{4}$, for $p>2$;
  6. (6)
    $A_{5}$, for $p=5$ or $p^{2m}-1\equiv 0\pmod{5}$;
  7. (7)
    Semidirect products of elementary abelian $p$-groups of order $p^{f}$ with cyclic groups of order $n$, with $f\mid m$, $n\mid(p^{f}-1)$ and $n\mid(p^{m}-1)$;
  8. (8)
    $\mathrm{PSL}(2,p^{f})$ and $\mathrm{PGL}(2,p^{f})$ with $f\mid m$.
The next lemma is verified using an elementary computation, so we omit the proof.
Lemma 3.8.
Let
$\sigma=\begin{pmatrix}-1&0\\ 0&1\end{pmatrix}\in PGL(2,q).$
Suppose $\eta\in PGL(2,q)$ satisfies the commutativity condition $\sigma\eta=\eta\sigma$. Then $\eta$ must be of the form
$\eta=\begin{pmatrix}a&0\\ 0&1\end{pmatrix}\quad\text{or}\quad\eta=\begin{pmatrix}0&1\\ b&0\end{pmatrix},$
for some $a,b\in\mathbb{F}_{q}$.
Another straightforward lemma turns out to be particularly useful in the $d=4$ case, and it may be of interest more generally. In fact, it is similar to, but simpler than, results in the case of six points as sketched in [6, p. 644-645].
Lemma 3.9.
Let $S=\{\alpha,-\alpha,1/\alpha,-1/\alpha\}$ be a subset of some field $k$, and assume #$\#S=4$ (equivalently, $\alpha^{4}\neq 1$). Then the group
$G_{S}:=\left\{f\in\operatorname{Aut}({\mathbb{P}}^{1})\;:\;f\;\text{permutes}\;S\right\}$
is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ and is generated by $\sigma\colon x\mapsto-x$ and $\tau\colon x\mapsto 1/x$, except in the following cases:
  1. (1)
    $S$ is the set of zeros of $x^{4}+1$;
  2. (2)
    $S$ is the set of zeros of $x^{4}+6x^{2}+1$;
  3. (3)
    $S$ is the set of zeros of $x^{4}-6x^{2}+1$;
  4. (4)
    $S$ is contained in the set of zeros of $x^{8}+14x^{4}+1$.
Proof.
Restriction to $S$ identifies $G_{S}$ with a subgroup of $S_{4}$. Under this identification, the group $\langle\sigma,\tau\rangle$ generated by $\sigma$ and $\tau$ corresponds to the normal subgroup of $S_{4}$ generated by (1 2)(3 4)1234(1\,2)(3\,4)( 1 2 ) ( 3 4 ) and (1 3)(2 4)1324(1\,3)(2\,4)( 1 3 ) ( 2 4 ), of order $4$. If $G_{S}$ is strictly larger than $\langle\sigma,\tau\rangle$, then either it contains a $2$-cycle, or a $3$-cycle. We now analyse these two possibilities.
If $G_{S}$ contains a $2$-cycle, after possibly multiplying by an element of $\langle\sigma,\tau\rangle$ we can assume it fixes $\alpha$ and one more element of $S$. In particular, considered as an $f\in\operatorname{Aut}({\mathbb{P}}^{1})$ we have $f(0)\neq 0$, which means $f$ has the form $f\colon x\mapsto\frac{ax+1}{cx+d}$. Moreover $f$ having order $2$ implies $d=-a$. Three possibilities remain:
[…]
The fixpoints show $f(x)=\alpha^{2}/x$ and therefore $-1/\alpha=f(1/\alpha)=\alpha^{3}$. This means that one is in case (1) of the lemma.
[…]
Now the fixpoints show $f(x)=\frac{(\alpha^{2}+1)x-2\alpha}{2\alpha x-\alpha^{2}-1}$. In this case the equality $f(-\alpha)=-1/\alpha$ is equivalent to $\alpha^{4}+6\alpha^{2}+1=0$, case (2) of the lemma.
[…]
Similar to the previous case one concludes here $f(x)=\frac{(\alpha^{2}-1)x+2\alpha}{2\alpha x+1-\alpha^{2}}$. From $f(-\alpha)=1/\alpha$ one deduces $\alpha^{4}-6\alpha^{2}+1=0$, case (3).
The possibility of a $3$-cycle is treated as follows. Conjugating with a suitable element in $\langle\sigma,\tau\rangle$, one may assume this map $f\in\operatorname{Aut}({\mathbb{P}}^{1})$ is determined by
$\alpha\mapsto\alpha,\;\;-\alpha\mapsto\frac{1}{\alpha}\mapsto-\frac{1}{\alpha}\mapsto-\alpha.$
Note that any $3$-cycle has two fixpoints in ${\mathbb{P}}^{1}$. For $f$, one is $\alpha$ and we claim that the other one is not $\infty$. Indeed, $f(\infty)=\infty$ means that $f\colon x\mapsto ax+b$ for certain $a\neq 0,b$. Considering how $f$ acts on ±$\pm\alpha$ and $1/\alpha$ leads to the conclusion $(\alpha^{2}-1)^{2}=0$, contradicting the assumption that #$\#S=4$.
Since $f$ does not fix $\infty$, it is of the form $f\colon x\mapsto\frac{ax+b}{x+d}$ with $b\neq ad$. In this case, the action of $f$ on $S$ shows that $a,b,d\in k$ are the unique solutions of
$\left(\begin{array}[]{ccc}\alpha&1&-\alpha\\ -\alpha^{2}&\alpha&-1\\ \alpha&\alpha^{2}&\alpha\\ 1&-\alpha&-\alpha^{2}\end{array}\right)\left(\begin{array}[]{c}a\\ b\\ d\end{array}\right)=\left(\begin{array}[]{c}\alpha^{2}\\ -\alpha\\ -1\\ -\alpha\end{array}\right).$
A straightforward calculation reveals that such solutions exist precisely in the two situations $\alpha^{4}+2\alpha^{3}+2\alpha^{2}-2\alpha+1=0$ and (reciprocal to the first one) $\alpha^{4}-2\alpha^{3}+2\alpha^{2}+2\alpha+1=0$. Since the product of these expressions is $\alpha^{8}+14\alpha^{4}+1$, this completes the proof. ∎
Now, we determine the automorphism group of ${\mathcal{C}}_{d}$ in the specific case $d=4$.
Theorem 3.10.
Let $k$ be an algebraically closed field $k$ of characteristic 2,5,7$\neq$ 2$\neq$ 5$\neq$ 7≠ 2 , ≠ 5 , ≠ 7. The automorphism group $G$ of the curve $y^{4}=\varphi_{4}(x)$ over $k$ has order $16$ and contains a normal subgroup $N\cong\mathbb{Z}/4\mathbb{Z}$ of order $4$ such that
$G/N\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}.$
Proof.
Since for $d=4$ we have $12d=p+$ 1 for $p=7$, Proposition 2.1 and our assumption that $\operatorname{char}(k)\neq 7$ imply that ${\mathcal{C}}_{4}$ has precisely $4$ total inflection points. Hence by Lemma 3.4 an exact sequence
$0\to\mathbb{Z}/4\mathbb{Z}\longrightarrow\operatorname{Aut}({\mathcal{C}}_{4})\longrightarrow\operatorname{Aut}({\mathbb{P}}^{1})$
exists, and the image of the rightmost arrow equals the group of fractional linear transformations preserving the set of zeros of the Chebyshev polynomial $\varphi_{4}(x)=x^{4}-4x^{2}+2$. In particular, this image can be seen as a subgroup of $S_{4}$. Let $\lambda^{4}=2$ and apply the transformation $x\mapsto\lambda x$. Under this change of variables, $\varphi_{4}(x)$ is converted into the separable polynomial
$2x^{4}-4\lambda^{2}x^{2}+2=2(x-\alpha)(x+\alpha)(x-\alpha^{-1})(x+\alpha^{-1})$
for some $\alpha\in k$.
Hence, writing $N$ the normal subgroup of $G=\operatorname{Aut}({\mathcal{C}}_{4})$ given above, by Lemma 3.9 and the observation $2\lambda^{2}\neq 0$, the group $G/N$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ except when $2\lambda^{2}=\pm 6$ or $x^{4}-2\lambda^{2}x^{2}+1$ divides $x^{8}+14x^{4}+1$. The first exception implies $836=4\lambda^{4}=$ 8, which does not occur by the assumption on $\operatorname{char}(k)$. Since $x^{8}+14x^{4}+1$ results in the remainder $2040\lambda^{2}x^{2}-$ 20 when divided by $x^{4}-2\lambda^{2}x^{2}+1$ and $\operatorname{char}(k)\neq 2$, the remaining exception is only possible for $\operatorname{char}(k)=5$.
This completes the proof. ∎
Remark 3.11.
In the case $\textrm{char}(k)=7$, the curve ${\mathcal{C}}_{4}\colon y^{4}=\varphi_{4}(x)$ contains $12$ total inflection points (see Proposition 2.1) and in this case Proposition 3.3 shows that the automorphism group is $\left(\mathbb{Z}/4\mathbb{Z}\times\mathbb{Z}/4\mathbb{Z}\right)\rtimes S_{3}$, of cardinality $96$.
In the remaining case $\operatorname{char}(k)=5$, the curve ${\mathcal{C}}_{4}$ has precisely $4$ total inflection points. With notations as in the statement of Theorem 3.10, the proof of Lemma 3.9 in fact shows that the group $G/N$ is the alternating group $A_{4}$. Since a nontrivial homomorphism $A_{4}\to\operatorname{Aut}(\mathbb{Z}/4\mathbb{Z})=\{\pm 1\}$ does not exist, one concludes in this situation
$\operatorname{Aut}({\mathcal{C}}_{4})\cong\mathbb{Z}/4\mathbb{Z}\rtimes A_{4},$
a group of order $48$. Explicitly ${\mathcal{C}}_{4}:~{}y^{4}=x^{4}-4x^{2}+2$ in characteristic $5$ admits the automorphism of order $3$ given by
$(x,y)\mapsto\left(\frac{-2\beta x+(1-2\beta^{2})}{(2\beta^{2}+2)x+\beta},\;\frac{\eta y}{(2\beta^{2}+2)x+\beta}\right),$
with $\beta^{4}-4\beta^{2}+2=0$, and $\eta=\beta^{3}+2\beta$. Note that ${\mathbb{F}}_{5}(\beta)={\mathbb{F}}_{5^{4}}$.
Lemma 3.12.
Assume $d>4$. Consider the curve $\mathcal{C}_{d}\colon y^{d}=\varphi_{d}(x)$ defined over a field $k$ of characteristic $p$, with $\gcd(p,d)=1$. If the curve admits an automorphism of order two given by
$(x,y)\mapsto\left(\frac{1}{bx},\frac{y}{cx}\right)$
for some $b,c\in k$, then $d$ must be even. Moreover, if $p>3$, then one of the following congruences holds:
$d\equiv 4\pmod{p},\quad\text{or}\quad d\equiv\frac{1}{2}\pmod{p}.$
Proof.
If $d$ is odd, then as automorphism of ${\mathbb{P}}^{2}$, a map of the given form sends $(0:0:1)\in{\mathcal{C}}_{d}$ to $(1:0:0)\not\in{\mathcal{C}}_{d}$. Hence the assumptions of the lemma imply that $d$ is even. In this case, given map defines an automorphism of ${\mathcal{C}}_{d}$ precisely when
$\varphi_{d}(x)=c^{d}x^{d}\varphi_{d}(\frac{1}{bx}).$
Comparing coefficients of the $x^{d}$-terms in these polynomials shows $c^{d}=(-1)^{d/2}/2$, and one obtains
$2\varphi_{d}(x)=(-1)^{d/2}x^{d}\varphi_{d}\left(\frac{1}{bx}\right).$
The coefficients of $x^{d-2}$ and $x^{2}$ (see the expression for them as presented on page 3), lead to
$8b^{2}=d.$
Similarly, the coefficients of $x^{d-4}$ yield (note that we assume $\operatorname{char}(k)=p>3$)
$8(d-3)=\frac{d(d^{2}-4)}{4!}\cdot\frac{1}{b^{4}}.$
Using both conditions yields
$(2d-1)(d-4)=0.$
This completes the proof.
Lemma 3.13.
Assume $14d=p^{r}+$ 1 and $d>4$. Denote by $G$ the automorphism group of the curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ (in characteristic $p$), and let the subgroup $N<G$ be given by
$N=\{(x,y)\mapsto(x,\zeta y)\mid\zeta^{d}=1\}.$
Then $N$ is normal in $G$ and $G/N$ is not isomorphic to $A_{4}$, $A_{5}$, $S_{4}$, $\mathrm{PSL}(2,p^{f})$, $\mathrm{PGL}(2,p^{f})$.
Proof.
Since $14d-$ 1 and $12d-$ 1 cannot both be powers of $p$, Proposition 2.1 shows that ${\mathcal{C}}_{d}$ contains precisely $d$ total inflection points. Hence Lemma 3.4 implies that $N$ is normal in $G$. Consider the element $\sigma\colon x\mapsto-x$ in $G/N$. From Lemma 3.12, if there exists an element $\eta\colon x\mapsto\frac{1}{bx}$ in $G/N$ for some constant $b$, then either $p=3$ or $d\equiv 4$ or $d\equiv\frac{1}{2}\pmod{p}$. We claim that $p\neq 3$. Indeed the condition $14d=3^{r}+$ 1 implies that $d$ is odd, contradicting Lemma 3.12.
Since $14d=p^{r}+$ 1, we have $d\not\equiv\frac{1}{2}\pmod{p}$. Hence $d\equiv 4\pmod{p}$, implying $16\equiv 4d\equiv 1\pmod{p}$, so $p=5$. However, this is impossible since the condition $14d=p^{r}+$ 1 implies $p\equiv-1\pmod{4}$.
The result now follows by observing that for any element of order two in the groups listed in the lemma, there exists another element of order two that commutes with it. Using Lemma 3.8, this does not hold for $\sigma\in G/N$. ∎
Remark 3.14.
For any $d$, there is an involution $x\mapsto-x$, hence the group $G/N$ cannot be an elementary abelian $p$-group.
Proposition 3.15.
Consider the curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ defined over a field $k$ of characteristic $p$, with $14d=p^{r}+$ 1 for some $r>0$. Then ${\mathcal{C}}_{d}$ admits the automorphism of order $3$ given by
$(x,y)\mapsto\left(\frac{2x+12}{2-x}\,,\,\frac{-4y}{2-x}\right).$
Proof.
Write $q=p^{r}$, then $14d=q+$ 1 and therefore $\sqrt{-1}\not\in{\mathbb{F}}_{q}$. This implies
$(-4)^{d}=(1+\sqrt{-1})^{q+1}=(1+\sqrt{-1})(1-\sqrt{-1})=2.$
It remains to show that the polynomial
$F(x):=\frac{(2-x)^{d}}{2}\varphi_{d}\left(\frac{2x+12}{2-x}\right)\in k[x]$
equals $\varphi_{d}(x)$. Recall that $\varphi_{d}(x)$ is the unique monic polynomial of degree $d$ such that
$\varphi_{d}\left(t+\frac{1}{t}\right)=t^{d}+\frac{1}{t^{d}}.$
It is straightforward to verify that
$t:=\frac{(\sqrt{x+2}+2)^{2}}{2-x}\in k(\sqrt{x+2})$
satisfies
$t+\frac{1}{t}=\frac{2x+12}{2-x}.$
Therefore, using $2-x=(\sqrt{2+x}+2)^{2}/t$ and $(2+\sqrt{x+2})(2-\sqrt{x+2})=2-x$ one concludes
$2F(x)=(2-x)^{d}\varphi_{d}(t+\frac{1}{t})=(\sqrt{x+2}+2)^{2d}+\frac{(2-x)^{2d}}{(\sqrt{x+2}+2)^{2d}}=(\sqrt{x+2}+2)^{2d}+(\sqrt{x+2}-2)^{2d}.$
Now substitute $x=s^{4}+\frac{1}{s^{4}}$, so that $x+2=(s^{2}+s^{-2})^{2}$ and $(\sqrt{x+2}\pm 2)^{2}=(s\pm\frac{1}{s})^{4}$. Then
$2F\left(s^{4}+\frac{1}{s^{4}}\right)=\left(s+\frac{1}{s}\right)^{4d}+\left(s-\frac{1}{s}\right)^{4d}=2\left(s^{d}+\frac{1}{s^{d}}\right),$
where it is used that $14d=q+$ 1. As a consequence $F(x)=\varphi_{d}(x)$, completing the proof. ∎
The next lemma will be used in the proof of Theorem 3.17 below.
Lemma 3.16.
Let $q\equiv 3\bmod 4$ be a power of a prime $p$ and assume $d=(q+1)/4>1$.
If $\mu=[x\mapsto(x-12c)/(cx+1)]\in\operatorname{Aut}({\mathbb{P}}^{1})$ permutes the zeros of $\varphi_{d}(x)$ and $c\neq 0$, then $c=\pm\frac{1}{2}$.
Proof.
As a polynomial over $\mathbb{Q}$, the splitting field of $\varphi_{d}(x)$ equals $\mathbb{Q}(\zeta+\zeta^{-1})$ with Galois group over $\mathbb{Q}$ equal to $({\mathbb{Z}}/4d{\mathbb{Z}})^{\times}/\{\pm 1\}$. The splitting field of $\varphi_{d}(x)$ over ${\mathbb{F}}_{p}$ therefore has Galois group isomorphic to the decomposition group at $p$ in $({\mathbb{Z}}/4d{\mathbb{Z}})^{\times}/\{\pm 1\}$, which is the subgroup generated by $\bar{p}$. The equality $14d=p^{r}+$ 1 shows that this subgroup has order $r$, hence the splitting field of $\varphi_{d}(x)$ over ${\mathbb{F}}_{p}$ equals ${\mathbb{F}}_{q}$. In particular, this means that $c\in{\mathbb{F}}_{q}^{\times}$. From the proof of Proposition 3.15, as $14d=q+$ 1, one obtains
$2\varphi_{d}(x)=\left(\sqrt{x+2}+2\right)^{2d}+\left(\sqrt{x+2}-2\right)^{2d}.$
Since $\mu$ permutes the zeros of $\varphi_{d}(x)$, this leads to
$\lambda\varphi_{d}(x)=\left(\sqrt{(1+2c)x+2-12c}+2\sqrt{cx+1}\right)^{2d}+\left(\sqrt{(1+2c)x+2-12c}-2\sqrt{cx+1}\right)^{2d},$
for some constant $\lambda\neq 0$. Squaring both sides and substituting $x=t+\frac{1}{t}$ (note that $14d=q+$ 1, so $\left((A+B)^{2d}+(A-B)^{2d}\right)^{2}=(A+B)^{q+1}+(A-B)^{q+1}+2(A^{2}-B^{2})^{2d}=2(A^{2})^{2d}+2(B^{2})^{2d}+2(A^{2}-B^{2})^{2d}$ and also $44^{2d}=2^{q+1}=$ 4 ), gives
$(\lambda^{2}/2)(t^{4d}+2t^{2d}+1)$ $=((2c+1)t^{2}+(2-12c)t+(1+2c))^{2d}$
$\quad+4(ct^{2}+t+c)^{2d}$
$\quad+((1-2c)t^{2}-(2+12c)t+(1-2c))^{2d}.$
Comparing degree $1$ coefficients in $t$ shows
$(1-6c)(1+2c)^{(q-1)/2}+2c^{(q-1)/2}-(1+6c)(1-2c)^{(q-1)/2}=0.$
Similarly, degree $2$ coefficients (note that $d\geq 4$ and $4d\equiv 1\bmod p$) yield
$(1+2c)^{2d}+(1-2c)^{2d}+4c^{2d}-c^{2d-2}-(1-6c)^{2}(1+2c)^{2d-2}-(1+6c)^{2}(1-2c)^{2d-2}=0,$
with exponents $22d=1+(q-1)/$ 2 and $22d-2=-1+(q-1)/$ 2.
In case $c\neq\pm\frac{1}{2}$, one has $c,1\pm 2c\in{\mathbb{F}}_{q}^{\times}$ and the $\deg(t)=1$ equality above leads to cases depending on which of $c,1\pm 2c$ are squares in ${\mathbb{F}}_{q}$:
  • (a)
    Suppose all three are squares, or all three are nonsquares. Then $01-6c+2-1-6c=$ 0, hence $p>3$ and $c=\frac{1}{6}$. Substituting this into the $\deg(t)=2$ equation leads to $p=7$ and $c=-1$ (and $d\equiv 2\bmod 7$). Hence one considers $\mu\colon x\mapsto(x-2)/(1-x)$ in characteristic $7$, which has order $6$ and fixpoints ±$\pm 3$ (the zeros of $\varphi_{2}(x)=x^{2}-2$). As $3$ is not a square in ${\mathbb{F}}_{q}$ for $q$ an odd power of $7$, we have $3^{2d}\equiv-3\bmod 7$ and hence here
    $(\lambda^{2}/2)(t^{4d}+2t^{2d}+1)=(t^{2}+1)^{2d}+4(t^{2}-t+1)^{2d}+4(t^{2}+t+1)^{2d}.$
    Since $d\equiv 2\bmod 7$, the coefficient of $t^{4}$ in the right-hand-side equals $4$, contradicting the assumption that $d>2$.
  • (b)
    In case only $1-2c$ is a nonsquare or is the only square, one obtains $01-6c+2+1+6c=$ 0, again a contradiction.
  • (c)
    In case only $c$ is a nonsquare and also when only $c$ is a square, we have $01-6c-2-1-6c=$ 0 hence $p>3$ and $c=-\frac{1}{6}$. This results using the $t^{2}$-coefficient in $p=7$, so $c=1$. We have the inverse of the map $\mu$ discussed in case (a) here, resulting in the same contradiction.
  • (d)
    If only $1+2c$ is a nonsquare, and also if only $1+2c$ is a square, then the $t$-coefficient does not give a further restriction. From the $t^{2}$-coefficient one concludes $012c^{2}+1=$ 0. This violates that $\mu\in\text{Aut}({\mathbb{P}}^{1})$, i.e., the matrix $\begin{pmatrix}1&-12c\\ c&1\end{pmatrix}$ is invertible.
This finishes the proof. ∎
Theorem 3.17.
Suppose that the characteristic of the field $k$ is $p\geq 3$. If $14d=q+$ 1, where $q$ is a power of $p$ and $d\geq 4$, then the automorphism group $G$ of the curve
${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$
has order $6d$ and contains a normal subgroup $N\cong\mathbb{Z}/d\mathbb{Z}$ such that
$G/N\cong S_{3}=D_{3}.$
Proof.
From Proposition 3.15, we know that $G/N<\operatorname{Aut}({\mathbb{P}}^{1})$ contains the subgroup
$\left\langle r:=[x\mapsto-x],\quad s:=[x\mapsto(2x+12)/(-x+2)]\mid r^{2}=s^{3}=1,\quad rsr=s^{-1}\right\rangle\cong D_{3}.$
By the classification of finite subgroups of $\operatorname{Aut}(\mathbb{P}^{1})$ (Theorem 3.7) and Lemma 3.13, we deduce that the possible structures for $G/N$ are as follows:
  • (I)
    Characteristic $p=3$, and $G/N$ is a semidirect product of an elementary abelian $3$-group $H$ and a cyclic group $K$.
  • (II)
    Characteristic $p>3$, and $G/N$ is a dihedral group.
Case (I): $p=3$.
Since $r$ has order two and, using Lemma 3.8 and the assumption $G/N\cong({\mathbb{Z}}/3{\mathbb{Z}})^{n}\rtimes{\mathbb{Z}}/m{\mathbb{Z}}$, no element in $G/N$ different from $\textit{id},r$ commutes with $r$. Hence $s\in H$ and $K=\langle r\rangle$. Since $H$ is abelian, any element
$\mu=[x\mapsto(ax+b)/(cx+d)]\in H$
commutes with $s$. This leads to the equality
$[x\mapsto((a+b)x+b)/((c+d)x+d)]=[x\mapsto(ax+b)/((a+c)x+b+d)],$
from which one obtains
$\mu=[x\mapsto x/(cx+1)]$
for some $c\neq 0$. By Lemma 3.16, this means $\mu=s^{\pm 1}$, completing the proof for $p=3$.
Case (II): $p>3$.
In this situation $G/N\cong D_{n}$, a dihedral group containing $D_{3}$, so $3|n$. Since $s$ has order exceeding $2$, it is in the cyclic group ${\mathbb{Z}}/n{\mathbb{Z}}<D_{n}$.
As any
$\mu=[x\mapsto(ax+b)/(cx+d)]\in G/N\cong D_{n}$
satisfies $\mu r\mu r=\textit{id}$, one concludes in case $\mu\neq r,\textit{id}$ that $a=d$. Assuming that moreover $\mu\in{\mathbb{Z}}/n{\mathbb{Z}}<D_{n}$, so that $\mu s=s\mu$, one derives $b=-12c$, so
$\mu=[x\mapsto(ax-12c)/(cx+a)].$
Taking $\mu$ of order $n$ (which is at least $3$), we conclude $a\neq 0$, hence
$\mu=[x\mapsto(x-12c)/(cx+1)].$
Lemma 3.16 now shows $\mu=s^{\pm 1}$. This means $n=3$, completing the proof. ∎
In many cases, Theorem 3.17 may be proven by a somewhat simpler argument:
Proposition 3.18.
Assume $\operatorname{char}(k)=p>3$, and let $d$ be an integer such that $3\nmid d$. If $14d=p^{r}+$ 1, then the automorphism group $G$ of the curve $y^{d}=\varphi_{d}(x)$ has order $6d$ and contains a normal subgroup $N$ of order $d$ such that $G/N\cong D_{3}$.
Proof.
Write $q:=p^{r}$. By the proof of Lemma 3.16, the condition $14d=q+$ 1 implies that $\varphi_{d}(x)$ splits completely in ${\mathbb{F}}_{q}[x]$. Moreover, the elements $x\mapsto-x$ and $x\mapsto(2x+12)/(2-x)$ (see Proposition 3.15) generate a subgroup $D_{3}\leq G/N$. As a consequence, Theorem 3.7 and Lemma 3.13 and Remark 3.14 imply that the quotient group $G/N\subset PGL(2,q)$ must be a dihedral group $D_{n}$ where $n$ divides $q\pm 1$ and $3|n$. Moreover, by Lemmas 3.8 and 3.12, one deduces that $n$ is odd.
The assumption that $3\nmid d$ leads to $n\nmid d$, implying that $n|(q+1)=4d$ cannot occur. Hence $n\mid(q-1)=4d-2$. Now take any $\sigma\in G/N\leq\operatorname{Aut}({\mathbb{P}}^{1})$ of order $n$. Its action on the zeros of $\varphi_{d}(x)$ is described by a product of pairwise disjoint $n$-cycles. Since $4d\equiv 2\bmod n$ and $n$ is odd, one obtains $d\equiv(n+1)/2\bmod n$, which means that $\sigma$ fixes $(n+1)/2$ of the zeros of $\varphi_{d}(x)$. From $3|n$ and the observation that $\sigma$ can have no more than $2$ fixed points, it follows that $n=3$. ∎
Based on calculations using Magma to find, for various small $d$ and characteristics $p\nmid 2d$, the elements in $\operatorname{Aut}({\mathbb{P}}^{1})$ restricting to permutations of the zeros of $\varphi_{d}(x)$, we are led to the following.
Expectation 3.19.
Assume $d>4$ and both $12d-$ 1 and $14d-$ 1 are not powers of the odd prime $p$. The curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ in characteristic $p$ has automorphism group isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/d\mathbb{Z}$.
As follows from Remark 3.11, the condition $d>4$ is necessary. Proposition 2.1 and Lemma 3.4 imply that an equivalent way to state the expectation is that under the conditions $d>4$ and $p\nmid 2d$ and $2d-1,4d-1\not\in p^{\mathbb{Z}}$, we expect
$\left\{\sigma\in\text{Aut}({\mathbb{P}}^{1}_{\overline{{\mathbb{F}}}_{p}})\;:\;\sigma\;\text{permutes the zeros of}\;\varphi(x)\right\}=\{x\mapsto\pm x\}.$
4. Nonisomorphic Maximal curves
Given a smooth, projective, and absolutely irreducible curve $X$ of genus $g$ over a finite field $\mathbb{F}_{q}$, the number of rational points satisfies the Hasse–Weil bound:
|#$|\#X(\mathbb{F}_{q})-(q+1)|\leq 2g\sqrt{q}.$
A curve is called maximal over $\mathbb{F}_{q}$ if it attains the upper bound, i.e.,
#$\#X(\mathbb{F}_{q})=q+1+2g\sqrt{q}.$
Maximal curves over finite fields play an important role in algebraic geometry and in coding theory. An obvious necessary condition for existence of a maximal curve over ${\mathbb{F}}_{q}$ is that $q$ is a square. A classical example is the Hermitian curve given by
$x^{q+1}+x^{q+1}+z^{q+1}=0,$
which is maximal over ${\mathbb{F}}_{q^{2}}$. A well known result following from J.T. Tate’s fundamental results on endomorphisms of abelian varieties over finite fields and attributed to J-P. Serre (see, e.g., [8, Prop. 6]), states that given curves ${\mathcal{C}},{\mathcal{D}}$ over ${\mathbb{F}}_{q^{2}}$ and a non-constant morphism ${\mathcal{C}}\to{\mathcal{D}}$ defined over ${\mathbb{F}}_{q^{2}}$, maximality of ${\mathcal{C}}$ over ${\mathbb{F}}_{q^{2}}$ implies maximality of ${\mathcal{D}}$ over ${\mathbb{F}}_{q^{2}}$. This is used frequently in what follows.
From [3, Theorem 4.4] it is known that the curve ${\mathcal{C}}_{d}\colon y^{d}=\varphi_{d}(x)$ is maximal over ${\mathbb{F}}_{q^{2}}$ if and only if $d$ divides $(q+1)/2$. In the special case $12d=q+$ 1, then Proposition 3.1 shows that the curve ${\mathcal{C}}_{d}$ is isomorphic to the Fermat curve ${\mathcal{F}}_{d}\colon x^{d}+y^{d}+z^{d}=0$ of degree $d=(q+1)/2$. Using Proposition 2.1, we know that if $d\neq(q+1)/2$, then the curve ${\mathcal{C}}_{d}$ has only $d$ total inflection points. This leads to a slightly different proof (but using the same ideas) of the following result presented in [1, Corollary 6.4]:
Proposition 4.1.
Let $q$ be a power of a prime $p>2$. For any divisor $n$ of $q+1$ satisfying
$3<n<\frac{q+1}{2},$
the smooth plane curve ${\mathcal{C}}_{n}$ is maximal over $\mathbb{F}_{q^{2}}$, and neither the Fermat curve ${\mathcal{F}}_{n}$ nor the Hurwitz curve ${\mathcal{H}}_{n-1}\colon xy^{n-1}+yz^{n-1}+zx^{n-1}=0$ is isomorphic to ${\mathcal{C}}_{n}$.
We now use our results on automorphism groups to obtain additional maximal curves (not necessarily plane) over $\mathbb{F}_{q^{2}}$ of the same genus that are not isomorphic.
Proposition 4.2.
Let $q$ be a power of a prime $p>3$ such that $\gcd(3,q+1)=1$. If $14n=q+$ 1, then for any divisor $m$ of $n$, the maximal curves ${\mathcal{C}}$ and $\mathcal{D}$ over $\mathbb{F}_{q^{2}}$ defined by
${\mathcal{C}}:y^{m}=\varphi_{n}(x)\quad\text{and}\quad\mathcal{D}:y^{m}=x^{n}+1$
have the same genus but are not isomorphic over $\mathbb{F}_{q^{2}}$.
Proof.
First we note that $n$ and $m$ divides $(q+1)/2$ and so we note that both of these curves are covered by the Hermitian curve and are therefore maximal over $\mathbb{F}_{q^{2}}$. The two curves evidently have the same genus. By [7, Theorem 1], the automorphism group of ${\mathcal{D}}$ does not contain any element of order $3$. However, from the proof of Proposition 3.15, one sees that ${\mathcal{C}}$ admits the following automorphism of order $3$:
$(x,y)\mapsto\left(\frac{2x+12}{-x+2},\frac{(-4)^{t}y}{(-x+2)^{t}}\right),$
where $n=tm$. This distinction implies that the two curves are not isomorphic. ∎
We conclude the paper by presenting another pair of non-isomorphic maximal curves with the same genus over $\mathbb{F}_{q^{2}}$.
Proposition 4.3.
Let $q$ be a power of a prime $p>3$ such that $q\equiv-1\pmod{8}$. Then, for any divisor $m\geq 3$ of $(q+1)/2$, the maximal curves $\mathcal{C}$ and $\mathcal{D}$ over $\mathbb{F}_{q^{2}}$, defined by
$\mathcal{C}:y^{m}=\varphi_{4}(x)\quad\text{and}\quad\mathcal{D}:y^{m}=x^{4}+1,$
have the same genus but are not isomorphic over $\mathbb{F}_{q^{2}}$.
Proof.
The conditions yield that both $4$ and $m$ divide $(q+1)/2$, hence ${\mathcal{C}}$ and ${\mathcal{D}}$ are covered by the Hermitian curve over ${\mathbb{F}}_{q^{2}}$. So both are maximal over ${\mathbb{F}}_{q^{2}}$, and clearly they have the same genus. If $m=4$, the result follows directly from Proposition 4.1. For the remaining cases, let $\zeta\in{\mathbb{F}}_{q^{2}}$ be a primitive $m$-th root of unity and consider in $\text{Aut}({\mathcal{C}})$ the subgroup $G$ of elements commuting with $(x,y)\mapsto(x,\zeta y)$. Following the arguments in the proof of Theorem 3.10, one shows that $G$ is isomorphic to either $\mathbb{Z}/m\mathbb{Z}\rtimes(\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z})$ or $\mathbb{Z}/m\mathbb{Z}\rtimes A_{4}$. Using [7, Theorem 1], one observes that in $\text{Aut}(\mathcal{D})$ the subgroup of all elements commuting with a given $\sigma$ of order $m$, is not of one of these types. Hence $\mathcal{C}$ and $\mathcal{D}$ are not isomorphic. ∎
Remark 4.4.
Note that for $m=2$, Proposition 4.3 would discuss the genus $1$ curves ${\mathcal{C}}\colon y^{2}=x^{4}-4x^{2}+2$ and ${\mathcal{D}}\colon y^{2}=x^{4}+1$. These curves have $j$-invariant $j({\mathcal{C}})=8000$ and $j({\mathcal{D}})=1728$. Since $8000-1728=2^{7}\cdot 7^{2}$, the unique characteristic such that both equations correspond to isomorphic and smooth elliptic curves, is $p=7$. One easily verifies that indeed the two elliptic curves are isomorphic (and maximal) over ${\mathbb{F}}_{49}$.